# The Computational Complexity of Counting Linear Regions in ReLU Neural Networks

### v2

@@ 2 Preliminaries
A ReLU neural network NN with $d\geq 0$ hidden layers is defined by $d+1$ affine transformations $T^{(i)}:\mathbb{R}^{n_{i-1}}\to\mathbb{R}^{n_{i}},x\mapsto A^{(i)}x+b^{(i)}$ for $i\in[d+1]$. We assume that $n_{0}=n$ and $n_{d+1}=1$. The ReLU network NN computes the CWPL function $f_{N}:\mathbb{R}^{n}\to\mathbb{R}$ with
@@
The matrices $A^{(i)}\in\mathbb{R}^{n_{i}\times n_{i-1}}$ are called the weights and the vectors $b^{(i)}\in\mathbb{R}^{n_{i}}$ are the biases of the $i$-th layer. We say the network has depth $d+1$ and size $s(N):=\sum_{i=1}^{d}n_{i}$. Equivalently, ReLU networks can also be represented as layered, directed, acyclic graphs where each dimension of each layer is represented by one vertex, called a neuron. Each neuron computes an affine transformation of the outputs of its predecessors, applies the ReLU function, and outputs the result. We denote the CPWL function mapping the network input to the output of a neuron $v$ by $f_{N,v}:\mathbb{R}^{n}\to\mathbb{R}$. If the reference to the ReLU network NN is clear, we abbreviate $f_{N,v}$ by $f_{v}$.
@@
Given a ReLU network NN, a vector $a\in\{0,1\}^{s(N)}$ is called an activation pattern of NN if there exists an input $x\in\mathbb{R}^{n}$ such that when NN receives $x$ as input, the $i$-th neuron in NN has positive output (is active) if $a_{i}=1$ and 0 if $a_{i}=0$. Given an activation pattern $a\in\{0,1\}^{s(N)}$, the network collapses to an affine function $f_{N}^{a}:\mathbb{R}^{n}\to\mathbb{R}$, and each neuron $i$ outputs an affine function $f^{a}_{N,i}:\mathbb{R}^{n}\to\mathbb{R}$ ($f^{a}_{N,i}$ is the zero function if $a_{i}=0$). Again, if the reference to the ReLU network NN is clear, we abbreviate $f^{a}_{N,i}$ by $f^{a}_{i}$.
@@ 3 Definitions of linear regions
Given a network NN and an activation pattern $a\in\{0,1\}^{s(N)}$ with support $I\subseteq[s(N)]$, the set $S_{N,a}=\{x\in\mathbb{R}^{n}:f_{i}^{a}(x)>0\;\text{ for all }i\in I,\;f_{i}^{a}(x)\leq 0\;\text{ for all }i\notin I\}$ is an activation region of NN. If the reference to the ReLU network is clear, we abbreviate $S_{N,a}$ by $S_{a}$.
@@
Given a ReLU network NN, a proper activation region of NN is a full-dimensional activation region of NN.
@@
Given a ReLU network NN and a polyhedral complex $\mathcal{P}$ that is compatible with $f_{N}$, a convex region of NN given $\mathcal{P}$ is a full-dimensional polyhedron $P\in\mathcal{P}$. The number of convex regions of NN is the minimum number of convex regions of any polyhedral complex $\mathcal{P}$ that is compatible with $f_{N}$.
Note that many different polyhedral complexes can attain this minimal number. Hence, in general, it is not possible to refer to a polyhedron $P$ as a ‘convex region of NN’ without specifying an associated polyhedral complex.
@@
Given a ReLU network NN, an open connected region of NN is an open, inclusion-wise maximal connected subset of $\mathbb{R}^{n}$ on which $f_{N}$ is affine.
@@
Given a ReLU network NN, a closed connected region of NN is a (closed) inclusion-wise maximal connected subset of $\mathbb{R}^{n}$ on which $f_{N}$ is affine.
@@
Figure 1: A ReLU network computing the function $f(x,y)=\max(-y,\min(0,-x))$. The closed connected regions (center left) and the activation regions (center right) are displayed. The slice $\{(x,0):x\geq 0\}$ is contained in the two closed connected regions with functions 0 (red) and y-y (blue). We have $R_{6}=R_{5}=R_{4}=3$, $R_{3}=4$, $R_{2}=6$ and $R_{1}=8$. In anti-clockwise direction starting from the region with value 0, the activation patterns are $010110$, $000000$, $001101$, $001001$, $101001$, $100000$, $110010$ and $110110$ (neurons are ordered from the upper left to the lower right).
@@
Given a ReLU network NN, an affine region of NN is an inclusion-wise maximal subset of $\mathbb{R}^{n}$ on which $f_{N}$ is affine.
For each definition, a linear region $S\subseteq\mathbb{R}^{n}$ of a ReLU network NN can be associated with an affine function $g:\mathbb{R}^{n}\to\mathbb{R}$ such that $f_{N}(x)=g(x)$ for all $x\in S$. The affine function $g$ is unique if $S$ is full-dimensional. We say that the function $g$ is computed or realized on $S$. If $g$ is the zero function, we call $S$ a zero region and a nonzero region otherwise. The following theorem is immediate from the definitions.
@@
Given a ReLU network NN, let $R_{1},R_{2},\dots,R_{6}$ denote the number of activation regions, proper activation regions, convex regions, open connected regions, closed connected regions and affine regions, respectively. Then: $R_{6}\leq R_{5}\leq R_{4}\leq R_{3}\leq R_{2}\leq R_{1}.$
@@
Input: A ReLU network NN.
Question: Does NN have strictly more than $K$ linear regions (according to a specified definition)?
@@
Input: A ReLU network NN.
Question: What is the number of linear regions of NN (according to a specified definition)?
@@ 4 Counting regions: one hidden layer
The idea of the proof is as follows. In a ReLU network NN with one hidden layer, each neuron corresponds to a hyperplane that divides the input space into two halfspaces. It is not guaranteed that each hyperplane also leads to a discontinuity of the gradient of $f_{N}$, since the functions of the neurons with the same hyperplane may add up to an affine function. The proof of Theorem˜4.1 shows that detecting whether a hyperplane of a neuron is canceled can be done in polynomial time. All hyperplanes of the network cancel if and only if the network computes an affine function and has thus only one linear region according to Definitions˜3, 4, 5 and 6. For Definitions˜1 and 2, 1-region-decision is trivial.
@@
Containment in $\mathsf{P}$ is easy for Definitions˜1 and 2, since an activation pattern $a\in\{0,1\}^{s(N)}$ of a ReLU network NN is a unique certificate for a (proper) activation region, which can be verified in polynomial time by computing the dimension of the set $S_{a}$, see Lemma˜A.5. More modifications are necessary to show $\mathsf{P}$ containment also for Definitions˜3 and 4.
@@ 5.1 Hardness of the decision version
Theorem˜5.2 can be proven using Lemma˜5.3 in two steps. First, we can extend the hardness result of Lemma˜5.3 from 1-region-decision to $K$-region-decision by adding a new function with $K$ linear regions, and second, we can increase the number of hidden layers of the resulting network from 22 to $L$ by adding L2L-2 additional hidden layers that compute the identity function.
@@
Question: Do the networks NN and $N^{\prime}$ compute the same function?
@@ 5.2 Hardness of exact and approximate counting
The reduction from $\mathsf{SAT}$ implies that approximating Linear region counting is intractable as well, in the following sense: We define an approximation algorithm achieving approximation ratio $\rho\leq 1$ as an algorithm that is guaranteed [^3] The algorithm may be probabilistic and return the correct answer with probability bounded away from $1/2$. to return, given a network NN as input, a number that is at least $\rho$ times the number of regions of NN. In fact, even though Lemma˜5.6 only holds for Definitions˜4 and 5, it is sufficient to prove the inapproximability result also for Definitions˜3 and 6.
@@
If a $\mathsf{SAT}$ formula has no satisfying assignment, the network produced by the reduction of Lemma˜5.6 will have exactly 11 linear region according to Definitions˜3, 4, 5 and 6. Otherwise, it will have at least $1+2^{n}$ linear regions according to Definitions˜3, 4, 5 and 6. If it was possible to achieve an approximation ratio larger than $(2^{n}+1)^{-1}$ in polynomial time we could decide if a $\mathsf{SAT}$ formula is satisfiable in polynomial time. This concludes the proof. ∎
@@
Given a $\mathsf{SAT}$ formula $\phi$, the network $N_{\phi}$ in the proof of Lemma˜5.3 has exactly one linear region if $\phi$ is unsatisfiable and at least 2 linear regions if $\phi$ is satisfiable. We proceed by showing that for a ReLU network NN with $R$ linear regions, the network computing the function $f^{(k)}:\mathbb{R}^{nk}\to\mathbb{R},\;f^{(k)}(x_{11},...,x_{1n},...,x_{k1},...,x_{kn})=\sum_{i=1}^{k}f_{N}(x_{i1},...,x_{in})$ has exactly $R^{k}$ linear regions. In words, $f^{(k)}$ is simply the sum of $k$ copies of $f_{N}$ each having as input a disjoint set of $n$ variables. Applying this construction to $N_{\phi}$ for appropriate values of $k$ gives the desired result. ∎
@@ 6 Counting regions using polynomial space
It is not hard to see that computing the number of activation regions and proper activation regions is possible in space that is polynomial in $\langle N\rangle$. Consider the following (informal) algorithm: Given a ReLU network NN, iterate over all $2^{s(N)}$ vectors in $\{0,1\}^{s(N)}$, and for each vector $a\in\{0,1\}^{s(N)}$, compute the dimension of $S_{a}$ in time polynomial in $\langle N\rangle$ (see Lemma˜A.5) to determine if $S_{a}$ is an activation region or a proper activation region, and increase a counter by one if this is the case.
@@
1:A ReLU network NN and a vector $(a_{1},\dots,a_{n},b)\in\mathbb{Q}^{n+1}$.
2:1 if $\sum_{i=1}^{n}a_{i}x_{i}+b$ is a function of an affine region of NN, else 0.
@@
1:A ReLU network NN.
2:Number of affine regions of NN.
@@ A.1 Details on the encoding size
The encoding size $\langle n\rangle$ of a nonnegative integer $n$ is $\langle n\rangle:=\lceil\log_{2}(n+1)\rceil$, the encoding size of a fraction $q=a/b$ with $a\in\mathbb{Z}$ and $b\in\mathbb{N}$ is $\langle q\rangle:=1+\langle|a|\rangle+\langle b\rangle$, and the encoding size of a matrix with rational entries $A=(a_{ij})_{i\in[n],\ j\in[m]}$ is $\langle A\rangle:=nm+\sum_{i\in[n],\ j\in[m]}\langle a_{ij}\rangle$. Since we are interested in the computational complexity, we restrict ourselves to ReLU networks with rational entries that can be represented with a finite number of bits. Then, a ReLU network NN of depth $d+1$ with $A^{(i)}\in\mathbb{Q}^{n_{i}\times n_{i-1}}$ and a vector $b^{(i)}\in\mathbb{Q}^{n_{i}}$ has encoding size $\langle N\rangle=\sum_{i=1}^{d+1}\langle A^{(i)}\rangle+\sum_{i=1}^{d+1}\langle b^{(i)}\rangle$. In particular, if $A_{\max}$ denotes the maximum encoding size of an entry in any $A^{(i)}$ and $b^{(i)}$, then $\langle N\rangle\leq(d+1)\cdot(\max\{n,n_{1},\dots,n_{d}\}+1)^{2}\cdot(1+A_{\max})=\text{poly}(n,s(N),A_{\max})$.
@@ A.2 Basic properties of linear regions
Given a ReLU network NN, the set of open connected regions of NN is equal to the unique set $\mathcal{S}$ with the minimal number of open connected subsets such that $\bigcup_{S\in\mathcal{S}}\overline{S}=\mathbb{R}^{n}$ and $f_{N}$ restricted to any $S\in\mathcal{S}$ is affine.
@@
By Zanotti [2025a, Lemma 3.2], the unique set $\mathcal{S}$ is then exactly the unique set of inclusion-maximal open subsets of $\mathbb{R}^{n}$ such that $f_{N}$ restricted to each inclusion-maximal subset is affine, which is by definition the set of open connected regions of NN. ∎
@@ A.5 Technical results
Given a ReLU network NN and an activation pattern $a\in\{0,1\}^{s(N)}$ corresponding to a proper activation region. Let $A_{\max}\in\mathbb{Z}$ be the maximum absolute value of any numerator or denominator in an entry of the matrices $A^{(1)},\dots,A^{(d+1)}$ and biases $b^{(1)},\dots,b^{(d+1)}$ and let $n_{\max}=\max\{n_{0},\dots,n_{d+1}\}$. Then, the encoding size of every coefficient of the affine function $f_{N}^{a}:\mathbb{R}^{n}\to\mathbb{R}$, $f_{N}^{a}(x)=\sum_{i=1}^{n}a_{i}x+b$ is bounded by
@@
Given an activation pattern $a\in\{0,1\}^{s(N)}$ of NN, we modify the matrices $A^{(1)},\dots,A^{(d+1)}$ and biases $b^{(1)},\dots,b^{(d+1)}$ by replacing the columns of matrices and bias entries corresponding to inactive neurons with 0 entries. Let $A^{(1),a},\dots,A^{(d+1),a}$ and biases $b^{(1),a},\dots,b^{(d+1),a}$ denote the modified matrices and biases. Then, a simple calculation shows that for all $x\in\mathbb{R}^{n}$, we have
@@
Given a ReLU network NN, an activation pattern $a\in\{0,1\}^{s(N)}$ and an index $i\in[s(N)]$ of a neuron, one can compute in time polynomial in $\langle N\rangle$ the (coefficients of the) affine function $f^{a}_{i}:\mathbb{R}^{n}\to\mathbb{R}$ which is computed at the output of the $i$-th neuron.
@@
Given a ReLU network NN and a vector $a\in\{0,1\}^{s(N)}$, one can compute the dimension of $S_{a}$ in time polynomial in $\langle N\rangle$.

### v3

@@ 2 Preliminaries
A ReLU neural network $N$ with $d\geq 0$ hidden layers is defined by $d+1$ affine transformations $T^{(i)}:\mathbb{R}^{n_{i-1}}\to\mathbb{R}^{n_{i}},x\mapsto A^{(i)}x+b^{(i)}$ for $i\in[d+1]$. We assume that $n_{0}=n$ and $n_{d+1}=1$. The ReLU network $N$ computes the CWPL function $f_{N}:\mathbb{R}^{n}\to\mathbb{R}$ with
@@
The matrices $A^{(i)}\in\mathbb{R}^{n_{i}\times n_{i-1}}$ are called the weights and the vectors $b^{(i)}\in\mathbb{R}^{n_{i}}$ are the biases of the $i$-th layer. We say the network has depth $d+1$ and size $s(N):=\sum_{i=1}^{d}n_{i}$. Equivalently, ReLU networks can also be represented as layered, directed, acyclic graphs where each dimension of each layer is represented by one vertex, called a neuron. Each neuron computes an affine transformation of the outputs of its predecessors, applies the ReLU function, and outputs the result. We denote the CPWL function mapping the network input to the output of a neuron $v$ by $f_{N,v}:\mathbb{R}^{n}\to\mathbb{R}$. If the reference to the ReLU network $N$ is clear, we abbreviate $f_{N,v}$ by $f_{v}$.
@@
Given a ReLU network $N$, a vector $a\in\{0,1\}^{s(N)}$ is called an activation pattern of $N$ if there exists an input $x\in\mathbb{R}^{n}$ such that when $N$ receives $x$ as input, the $i$-th neuron in $N$ has positive output (is active) if $a_{i}=1$ and 0 if $a_{i}=0$. Given an activation pattern $a\in\{0,1\}^{s(N)}$, the network collapses to an affine function $f_{N}^{a}:\mathbb{R}^{n}\to\mathbb{R}$, and each neuron $i$ outputs an affine function $f^{a}_{N,i}:\mathbb{R}^{n}\to\mathbb{R}$ ($f^{a}_{N,i}$ is the zero function if $a_{i}=0$). Again, if the reference to the ReLU network $N$ is clear, we abbreviate $f^{a}_{N,i}$ by $f^{a}_{i}$.
@@ 3 Definitions of linear regions
Given a network $N$ and an activation pattern $a\in\{0,1\}^{s(N)}$ with support $I\subseteq[s(N)]$, the set $S_{N,a}=\{x\in\mathbb{R}^{n}:f_{i}^{a}(x)>0\;\text{ for all }i\in I,\;f_{i}^{a}(x)\leq 0\;\text{ for all }i\notin I\}$ is an activation region of $N$. If the reference to the ReLU network is clear, we abbreviate $S_{N,a}$ by $S_{a}$.
@@
Given a ReLU network $N$, a proper activation region of $N$ is a full-dimensional activation region of $N$.
@@
Given a ReLU network $N$ and a polyhedral complex $\mathcal{P}$ that is compatible with $f_{N}$, a convex region of $N$ given $\mathcal{P}$ is a full-dimensional polyhedron $P\in\mathcal{P}$. The number of convex regions of $N$ is the minimum number of convex regions of any polyhedral complex $\mathcal{P}$ that is compatible with $f_{N}$.
Note that many different polyhedral complexes can attain this minimal number. Hence, in general, it is not possible to refer to a polyhedron $P$ as a ‘convex region of $N$’ without specifying an associated polyhedral complex.
@@
Given a ReLU network $N$, an open connected region of $N$ is an open, inclusion-wise maximal connected subset of $\mathbb{R}^{n}$ on which $f_{N}$ is affine.
@@
Given a ReLU network $N$, a closed connected region of $N$ is a (closed) inclusion-wise maximal connected subset of $\mathbb{R}^{n}$ on which $f_{N}$ is affine.
@@
Figure 1: A ReLU network computing the function $f(x,y)=\max(-y,\min(0,-x))$. The closed connected regions (center left) and the activation regions (center right) are displayed. The slice $\{(x,0):x\geq 0\}$ is contained in the two closed connected regions with functions 0 (red) and y-y (blue). We have $R_{6}=R_{5}=R_{4}=3$, $R_{3}=4$, $R_{2}=5$ and $R_{1}=7$. In anti-clockwise direction starting from the region with value 0, the activation patterns are $010110$, $000000$, $001001$, $101001$, $100000$, $110010$ and $110110$ (neurons are ordered from the upper left to the lower right).
@@
Given a ReLU network $N$, an affine region of $N$ is an inclusion-wise maximal subset of $\mathbb{R}^{n}$ on which $f_{N}$ is affine.
For each definition, a linear region $S\subseteq\mathbb{R}^{n}$ of a ReLU network $N$ can be associated with an affine function $g:\mathbb{R}^{n}\to\mathbb{R}$ such that $f_{N}(x)=g(x)$ for all $x\in S$. The affine function $g$ is unique if $S$ is full-dimensional. We say that the function $g$ is computed or realized on $S$. If $g$ is the zero function, we call $S$ a zero region and a nonzero region otherwise. The following theorem is immediate from the definitions.
@@
Given a ReLU network $N$, let $R_{1},R_{2},\dots,R_{6}$ denote the number of activation regions, proper activation regions, convex regions, open connected regions, closed connected regions and affine regions, respectively. Then: $R_{6}\leq R_{5}\leq R_{4}\leq R_{3}\leq R_{2}\leq R_{1}.$
@@
Input: A ReLU network $N$.
Question: Does $N$ have strictly more than $K$ linear regions (according to a specified definition)?
@@
Input: A ReLU network $N$.
Question: What is the number of linear regions of $N$ (according to a specified definition)?
@@ 4 Counting regions: one hidden layer
The idea of the proof is as follows. In a ReLU network $N$ with one hidden layer, each neuron corresponds to a hyperplane that divides the input space into two halfspaces. It is not guaranteed that each hyperplane also leads to a discontinuity of the gradient of $f_{N}$, since the functions of the neurons with the same hyperplane may add up to an affine function. The proof of Theorem˜4.1 shows that detecting whether a hyperplane of a neuron is canceled can be done in polynomial time. All hyperplanes of the network cancel if and only if the network computes an affine function and has thus only one linear region according to Definitions˜3, 4, 5 and 6. For Definitions˜1 and 2, 1-region-decision is trivial.
@@
Containment in $\mathsf{P}$ is easy for Definitions˜1 and 2, since an activation pattern $a\in\{0,1\}^{s(N)}$ of a ReLU network $N$ is a unique certificate for a (proper) activation region, which can be verified in polynomial time by computing the dimension of the set $S_{a}$, see Lemma˜A.5. More modifications are necessary to show $\mathsf{P}$ containment also for Definitions˜3 and 4.
@@ 5.1 Hardness of the decision version
Theorem˜5.2 can be proven using Lemma˜5.3 in two steps. First, we can extend the hardness result of Lemma˜5.3 from 1-region-decision to $K$-region-decision by adding a new function with $K$ linear regions, and second, we can increase the number of hidden layers of the resulting network from $2$ to $L$ by adding L2L-2 additional hidden layers that compute the identity function.
@@
Question: Do the networks $N$ and $N^{\prime}$ compute the same function?
@@ 5.2 Hardness of exact and approximate counting
The reduction from $\mathsf{SAT}$ implies that approximating Linear region counting is intractable as well, in the following sense: We define an approximation algorithm achieving approximation ratio $\rho\leq 1$ as an algorithm that is guaranteed [^3] The algorithm may be probabilistic and return the correct answer with probability bounded away from $1/2$. to return, given a network $N$ as input, a number that is at least $\rho$ times the number of regions of $N$. In fact, even though Lemma˜5.6 only holds for Definitions˜4 and 5, it is sufficient to prove the inapproximability result also for Definitions˜3 and 6.
@@
If a $\mathsf{SAT}$ formula has no satisfying assignment, the network produced by the reduction of Lemma˜5.6 will have exactly $1$ linear region according to Definitions˜3, 4, 5 and 6. Otherwise, it will have at least $1+2^{n}$ linear regions according to Definitions˜3, 4, 5 and 6. If it was possible to achieve an approximation ratio larger than $(2^{n}+1)^{-1}$ in polynomial time we could decide if a $\mathsf{SAT}$ formula is satisfiable in polynomial time. This concludes the proof. ∎
@@
Given a $\mathsf{SAT}$ formula $\phi$, the network $N_{\phi}$ in the proof of Lemma˜5.3 has exactly one linear region if $\phi$ is unsatisfiable and at least 2 linear regions if $\phi$ is satisfiable. We proceed by showing that for a ReLU network $N$ with $R$ linear regions, the network computing the function $f^{(k)}:\mathbb{R}^{nk}\to\mathbb{R},\;f^{(k)}(x_{11},...,x_{1n},...,x_{k1},...,x_{kn})=\sum_{i=1}^{k}f_{N}(x_{i1},...,x_{in})$ has exactly $R^{k}$ linear regions. In words, $f^{(k)}$ is simply the sum of $k$ copies of $f_{N}$ each having as input a disjoint set of $n$ variables. Applying this construction to $N_{\phi}$ for appropriate values of $k$ gives the desired result. ∎
@@ 6 Counting regions using polynomial space
It is not hard to see that computing the number of activation regions and proper activation regions is possible in space that is polynomial in $\langle N\rangle$. Consider the following (informal) algorithm: Given a ReLU network $N$, iterate over all $2^{s(N)}$ vectors in $\{0,1\}^{s(N)}$, and for each vector $a\in\{0,1\}^{s(N)}$, compute the dimension of $S_{a}$ in time polynomial in $\langle N\rangle$ (see Lemma˜A.5) to determine if $S_{a}$ is an activation region or a proper activation region, and increase a counter by one if this is the case.
@@
1:A ReLU network $N$ and a vector $(a_{1},\dots,a_{n},b)\in\mathbb{Q}^{n+1}$.
2:1 if $\sum_{i=1}^{n}a_{i}x_{i}+b$ is a function of an affine region of $N$, else 0.
@@
1:A ReLU network $N$.
2:Number of affine regions of $N$.
@@ A.1 Details on the encoding size
The encoding size $\langle n\rangle$ of a nonnegative integer $n$ is $\langle n\rangle:=\lceil\log_{2}(n+1)\rceil$, the encoding size of a fraction $q=a/b$ with $a\in\mathbb{Z}$ and $b\in\mathbb{N}$ is $\langle q\rangle:=1+\langle|a|\rangle+\langle b\rangle$, and the encoding size of a matrix with rational entries $A=(a_{ij})_{i\in[n],\ j\in[m]}$ is $\langle A\rangle:=nm+\sum_{i\in[n],\ j\in[m]}\langle a_{ij}\rangle$. Since we are interested in the computational complexity, we restrict ourselves to ReLU networks with rational entries that can be represented with a finite number of bits. Then, a ReLU network $N$ of depth $d+1$ with $A^{(i)}\in\mathbb{Q}^{n_{i}\times n_{i-1}}$ and a vector $b^{(i)}\in\mathbb{Q}^{n_{i}}$ has encoding size $\langle N\rangle=\sum_{i=1}^{d+1}\langle A^{(i)}\rangle+\sum_{i=1}^{d+1}\langle b^{(i)}\rangle$. In particular, if $A_{\max}$ denotes the maximum encoding size of an entry in any $A^{(i)}$ and $b^{(i)}$, then $\langle N\rangle\leq(d+1)\cdot(\max\{n,n_{1},\dots,n_{d}\}+1)^{2}\cdot(1+A_{\max})=\text{poly}(n,s(N),A_{\max})$.
@@ A.2 Basic properties of linear regions
Given a ReLU network $N$, the set of open connected regions of $N$ is equal to the unique set $\mathcal{S}$ with the minimal number of open connected subsets such that $\bigcup_{S\in\mathcal{S}}\overline{S}=\mathbb{R}^{n}$ and $f_{N}$ restricted to any $S\in\mathcal{S}$ is affine.
@@
By Zanotti [2025a, Lemma 3.2], the unique set $\mathcal{S}$ is then exactly the unique set of inclusion-maximal open subsets of $\mathbb{R}^{n}$ such that $f_{N}$ restricted to each inclusion-maximal subset is affine, which is by definition the set of open connected regions of $N$. ∎
@@ A.5 Technical results
Given a ReLU network $N$ and an activation pattern $a\in\{0,1\}^{s(N)}$ corresponding to a proper activation region. Let $A_{\max}\in\mathbb{Z}$ be the maximum absolute value of any numerator or denominator in an entry of the matrices $A^{(1)},\dots,A^{(d+1)}$ and biases $b^{(1)},\dots,b^{(d+1)}$ and let $n_{\max}=\max\{n_{0},\dots,n_{d+1}\}$. Then, the encoding size of every coefficient of the affine function $f_{N}^{a}:\mathbb{R}^{n}\to\mathbb{R}$, $f_{N}^{a}(x)=\sum_{i=1}^{n}a_{i}x+b$ is bounded by
@@
Given an activation pattern $a\in\{0,1\}^{s(N)}$ of $N$, we modify the matrices $A^{(1)},\dots,A^{(d+1)}$ and biases $b^{(1)},\dots,b^{(d+1)}$ by replacing the columns of matrices and bias entries corresponding to inactive neurons with 0 entries. Let $A^{(1),a},\dots,A^{(d+1),a}$ and biases $b^{(1),a},\dots,b^{(d+1),a}$ denote the modified matrices and biases. Then, a simple calculation shows that for all $x\in\mathbb{R}^{n}$, we have
@@
Given a ReLU network $N$, an activation pattern $a\in\{0,1\}^{s(N)}$ and an index $i\in[s(N)]$ of a neuron, one can compute in time polynomial in $\langle N\rangle$ the (coefficients of the) affine function $f^{a}_{i}:\mathbb{R}^{n}\to\mathbb{R}$ which is computed at the output of the $i$-th neuron.
@@
Given a ReLU network $N$ and a vector $a\in\{0,1\}^{s(N)}$, one can compute the dimension of $S_{a}$ in time polynomial in $\langle N\rangle$.